OpenStudy (anonymous):

A hot air balloon holds 4,328 cubic meters of helium. The density of helium is 0.1785 kilograms per cubic meter. How many kilograms of helium does the balloon contain, rounded to the nearest tenth of a kilogram? A. 3,747.9 kg B. 3,875.3 kg C. 772.5 kg D. 4,383.2 kg

OpenStudy (anonymous):

@phi @magbak @mathslover @KingGeorge @ajprincess

mathslover (mathslover):

Use the formula of density here : \(\rho = \cfrac{mass}{volume}\) where \(\rho\) ( pronounced as rho) is the density of the substance (here substance is Helium) Given to you that density of helium in the balloon : \(\rho _ {He} = 0.1785 kg m^{-3}\) And given that volume of Helium gas in the balloon = 4,328 \(m^3\)

mathslover (mathslover):

So, just put the values : \(\rho _{He} = \cfrac{m_{He}}{V_{He}}\)

OpenStudy (anonymous):

I still dont understand...

mathslover (mathslover):

What is given to you in the question?

OpenStudy (anonymous):

The amount of helium in the balloon and the density.

mathslover (mathslover):

Right. So, it is actually volume of helium gas and the density Since, density = mass/ volume So, mass = density * volume

mathslover (mathslover):

Just put the given values.

OpenStudy (anonymous):

4,328/0.1785 = 24,246...

mathslover (mathslover):

It is multiplication, not division .

mathslover (mathslover):

That is Mass = Volume * Density = 4328 * 0.1785

OpenStudy (anonymous):

Got it, Thank you!

mathslover (mathslover):

You're very welcome :)

OpenStudy (shaw74):

the answer is c

OpenStudy (anonymous):

4328 * 0.1785 = 772.548 So C is correct!